The cost of equipment purchased by Waterway, Inc., on June 1, 2017, is $95,230. It is estimated that the machine will have a $5,350 salvage value at the end of its service life. Its service life is estimated at 7 years, its total working hours are estimated at 44,940, and its total production is estimated at 561,750 units. During 2017, the machine was operated 6,420 hours and produced 58,850 units. During 2018, the machine was operated 5,885 hours and produced 51,360 units.

Compute depreciation expense on the machine for the year ending December 31, 2017, and the year ending December 31, 2018, using the following methods. (Round depreciation per unit to 2 decimal places, e.g. 15.25 and final answers to 0 decimal places, e.g. 45,892.)

2017

2018

(a) Straight-line $

$

(b) Units-of-output $

$

(c) Working hours $

$

(d) Sum-of-the-years'-digits $

$

(e) Double-declining-balance (twice the straight-line rate) $

$
Answer #1

Compute depreciation expense on the machine for the year ending December 31, 2017, and the year ending December 31, 2018, using the following methods. (Round depreciation per unit to 2 decimal places, e.g. 15.25 and final answers to 0 decimal places, e.g. 45,892.)

2017

2018

(a) Straight-line $89880/7*7/12 = 7490

$12840

(b) Units-of-output 58850*.16 = 9416

$8218

(c) Working hours $12840

$11770

(d) Sum-of-the-years'-digits $89880*7/28*7/12 = 13108

$89880*7/28*5/12+89880*6/28*7/12 = 20597

(e) Double-declining-balance (twice the straight-line rate) $95230*28.57%*7/12 = 15871 $11336+11337 = 22673